# Dark energy and matter interacting scenario can relieve $H_{0}$ and $S_{8}$ tensions

### v2

@@ Abstract
In the CBS+$H_{0}$ case, we have $H_{0}=72.3\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.817\pm 0.010$. So in this case the $H_{0}$ and $S_{8}$ tensions are further relieved, and we can see that the $H_{0}$ tension is greatly relieved to $0.47\sigma$. What’s more, in this case $\Delta{\rm AIC}=-2.07$, indicating that the $\tilde{\Lambda}$ model is favored by the data over $\Lambda$CDM. However, the $\tilde{\Lambda}$ model is not favored by the CBS+$S_{8}$ data, in which though the $S_{8}$ tension is slightly improved (to $2.43\sigma$), the $H_{0}$ tension is slightly relieved (to $1.96\sigma$), and the fit gives $\Delta{\rm AIC}=5.02$. Consequently, incorporating $H_{0}$ and $S_{8}$ priors into the data combination effectively mitigates the tensions. Therefore, employing the CBS+$H_{0}$+$S_{8}$ data set fosters an optimally synthesized scenario. In this scenario, we obtain $H_{0}=71.9\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.8074\pm 0.00998$, consequently reducing the $H_{0}$ and $S_{8}$ tensions to $0.72\sigma$ and $2.11\sigma$, respectively. Furthermore, under these conditions, the model exhibits the highest degree of concordance with the data, as indicated by $\Delta{\rm AIC}=-2.29$.
@@
We compare our findings with those derived from the simplified $\tilde{\Lambda}$ model Gao et al. (2021), wherein it is postulated that $\delta_{G}^{M}$ and $\delta_{G}^{R}$ are identical in Eqs. (5) and (7), i.e., $\delta_{G}^{M}=\delta_{G}^{R}=\delta_{G}$, and $\delta_{\Lambda}=\delta_{G}(\Omega_{\rm m}+\Omega_{\rm r})/\Omega_{\Lambda}$. Employing the CBS data set, the $H_{0}$ tension can be relieved to $3.59\sigma$ (with the best fit $H_{0}=67.71{\rm km~{}s^{-1}~{}Mpc^{-1}}$) and in this case the $S_{8}$ tension is exacerbated (with the best fit $S_{8}=0.8252$). Additionally, within the simplified $\tilde{\Lambda}$ model constrained by the CBS + $H_{0}$ data set, we have the best-fit values $\sigma_{8}=0.8720$ and $S_{8}=0.8310$, and we can see that the $\sigma_{8}$ and $S_{8}$ tensions significantly increase while alleviating the $H_{0}$ tension.
@@
Additionally, we extend our comparison to the $\Lambda(t)$CDM model, an interacting vacuum energy model that posits an energy exchange between vacuum energy and cold dark matter as detailed by specific equations $\dot{\rho}_{\Lambda}=\beta H\rho_{c}$ and $\dot{\rho}_{c}+3H\rho_{c}=-\beta H\rho_{c}$ (here the dimensionless parameter $\beta$ describes the interaction strength). In Ref. Guo et al. (2019), the same CBS+$H_{0}$ data set is used to constrain $\Lambda(t)$CDM model to obtain $H_{0}=69.36~{}{\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $\sigma_{8}=0.844$. In contrast, the current findings, as illustrated in Table 2, explicitly show that the $\tilde{\Lambda}$ model exhibits distinct advantages relative to the $\Lambda(t)$CDM model.
@@
Conclusion.—In this work, we proposed a new cosmological model in which the vacuum energy interacts with matter and radiation, which is considered to originate from the asymptotic safety and particle production of gravitational quantum field theory. We tested this model using the current cosmological observations and discussed its capability of relieving the $H_{0}$ and $S_{8}$ tensions.
@@
To elucidate the mechanisms by which the ${\tilde{\Lambda}}$ model ameliorates tensions, we analyzed the differences $\delta\Omega_{{\rm m,r},\Lambda}(z)$ between the ${\tilde{\Lambda}}$ and ${\Lambda}$CDM models. The alleviation of the $H_{0}$ tension is attributed to the ${\tilde{\Lambda}}$ model’s enhanced dark energy proportion and reduced matter fraction at low redshifts relative to the ${\Lambda}$CDM model, stemming from the conversion of matter into dark energy. The findings of $\delta_{G}^{M,R}<0$ and $\delta_{\Lambda}<0$ corroborate the $\tilde{\Lambda}$ hypothesis in which in the late universe radiation and matter decay into dark energy Xue (2022, 2023c).
@@
We find that this cosmological model can significantly relieve the $H_{0}$ tension, and at the same time, it can also slightly reduce the $S_{8}$ tension, which cannot be easily observed in other cosmological models. When using the CBS data to constrain the model, we find that the $H_{0}$ tension is relieved to $2.87\sigma$, and meanwhile, the $S_{8}$ tension is also improved to $2.77\sigma$. But in this case, the $S_{8}$ tension is only slightly reduced, and actually, the model is not favored by the CBS data (since $\Delta{\rm AIC}=3.79$).
@@
When the $H_{0}$ and $S_{8}$ data are added into the data combination, the situation is significantly improved. In the CBS+$H_{0}$ case, we obtain the result of $H_{0}=72.3\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$, indicating that the $H_{0}$ tension is relieved to $0.47\sigma$, and in this case the model is favored over $\Lambda$CDM ($\Delta{\rm AIC}=-2.07$). In the CBS+$H_{0}$+$S_{8}$ case, we get a synthetically best situation, in which we have $H_{0}=71.9\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.8074\pm 0.0098$, thus the $H_{0}$ and $S_{8}$ tensions are relived to $0.72\sigma$ and $2.11\sigma$, respectively. In this case, the model is most favored by the data ($\Delta{\rm AIC}=-2.29$).
@@
Therefore, we find that such a cosmological model can greatly relieve the $H_{0}$ tension, and at the same time, it can also alleviate the $S_{8}$ tension.
@@
Undoubtedly, this model requires further in-depth research in many aspects. Such an interaction between vacuum energy and matter is likely to bring many additional observational effects. For example, we are not sure whether this model will lead to a significant integrated Sachs-Wolfe effect. Further research on this issue requires further related theoretical study and analysis using full CMB angular power spectrum data. Another example is that this additional interaction may also lead to modifications in the middle-scale clustering patterns in large-scale structures, which will potentially affect the clustering strength at various scales. Therefore, such modifications may also affect the mass function and profile of dark matter halos, the statistical properties of galaxy clusters, the alignments of galaxies, the structure of the cosmic web, and so on. The study of these effects is a complicated issue and requires the use of N-body simulations, and also relies on assumptions about the nature of dark matter (such as cold dark matter or fuzzy dark matter). All these aspects deserve further in-depth discussion.

### v3

@@ Abstract
In the CBS+$H_{0}$ case, we have $H_{0}=72.3\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.817\pm 0.010$. Therefore, the $H_{0}$ and $S_{8}$ tensions are further relieved, and the $H_{0}$ tension is greatly relieved to $0.47\sigma$. Furthermore, in this case, $\Delta{\rm AIC}=-2.07$, indicating that the $\tilde{\Lambda}$ model is favored by the data over $\Lambda$CDM. However, the $\tilde{\Lambda}$ model is not favored by the CBS+$S_{8}$ data, in which though the $S_{8}$ tension is slightly improved (to $2.43\sigma$), the $H_{0}$ tension is slightly relieved (to $1.96\sigma$), and the fit gives $\Delta{\rm AIC}=5.02$. Consequently, incorporating $H_{0}$ and $S_{8}$ priors into the data combination effectively mitigates the tensions. Therefore, employing the CBS+$H_{0}$+$S_{8}$ dataset fosters an optimally synthesized scenario, for which we obtain $H_{0}=71.9\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.8074\pm 0.00998$, consequently reducing the $H_{0}$ and $S_{8}$ tensions to $0.72\sigma$ and $2.11\sigma$, respectively. Furthermore, under these conditions, the model exhibits the highest degree of concordance with the data, as indicated by $\Delta{\rm AIC}=-2.29$.
@@
We compare our findings with those derived from the simplified $\tilde{\Lambda}$ model Gao et al. (2021), wherein it is postulated that $\delta_{G}^{M}$ and $\delta_{G}^{R}$ are identical in Eqs. (5) and (7), that is, $\delta_{G}^{M}=\delta_{G}^{R}=\delta_{G}$ and $\delta_{\Lambda}=\delta_{G}(\Omega_{\rm m}+\Omega_{\rm r})/\Omega_{\Lambda}$. Employing the CBS dataset, the $H_{0}$ tension can be relieved to $3.59\sigma$ (with the best fit $H_{0}=67.71~{}{\rm km~{}s^{-1}~{}Mpc^{-1}}$), and in this case, the $S_{8}$ tension is exacerbated (with the best fit $S_{8}=0.8252$). Additionally, within the simplified $\tilde{\Lambda}$ model constrained by the CBS + $H_{0}$ dataset, we obtain the best-fit values $\sigma_{8}=0.8720$ and $S_{8}=0.8310$, and the $\sigma_{8}$ and $S_{8}$ tensions significantly increase while alleviating the $H_{0}$ tension.
@@
We extend our comparison to the $\Lambda(t)$CDM model, an interacting vacuum energy model that posits an energy exchange between vacuum energy and cold dark matter as detailed by the specific equations $\dot{\rho}_{\Lambda}=\beta H\rho_{c}$ and $\dot{\rho}_{c}+3H\rho_{c}=-\beta H\rho_{c}$ (here, the dimensionless parameter $\beta$ describes the interaction strength). In Ref. Guo et al. (2019), the same CBS+$H_{0}$ dataset is used to constrain the $\Lambda(t)$CDM model to obtain $H_{0}=69.36~{}{\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $\sigma_{8}=0.844$. In contrast, as illustrated in Table 2, the current findings explicitly show that the $\tilde{\Lambda}$ model exhibits distinct advantages over the $\Lambda(t)$CDM model.
@@
Conclusion.—In this study, we propose a new cosmological model, in which the vacuum energy interacts with matter and radiation, which is considered to originate from the asymptotic safety and particle production of gravitational quantum field theory. We test this model using the current cosmological observations and discuss its capability for relieving the $H_{0}$ and $S_{8}$ tensions.
@@
To elucidate the mechanisms by which the ${\tilde{\Lambda}}$ model relieves tensions, we analyze the differences $\delta\Omega_{{\rm m,r},\Lambda}(z)$ between the ${\tilde{\Lambda}}$ and ${\Lambda}$CDM models. The alleviation of the $H_{0}$ tension is attributed to the ${\tilde{\Lambda}}$ model’s enhanced dark energy proportion and reduced matter fraction at low redshifts relative to the ${\Lambda}$CDM model, which stem from the conversion of matter into dark energy. The findings of $\delta_{G}^{M,R}<0$ and $\delta_{\Lambda}<0$ corroborate the $\tilde{\Lambda}$ hypothesis in which in the late universe radiation and matter decay into dark energy Xue (2022, 2023c).
@@
We find that this cosmological model can significantly relieve the $H_{0}$ tension and simultaneously slightly reduce the $S_{8}$ tension, which cannot be easily observed in other cosmological models. When using the CBS data to constrain the model, we find that the $H_{0}$ tension is relieved to $2.87\sigma$, and meanwhile, the $S_{8}$ tension is also improved to $2.77\sigma$. However, in this case, the $S_{8}$ tension is only slightly reduced, and the model is not favored by the CBS data (because $\Delta{\rm AIC}=3.79$).
@@
When the $H_{0}$ and $S_{8}$ data are added to the data combination, the situation is significantly improved. In the CBS+$H_{0}$ case, we obtain the result $H_{0}=72.3\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$, indicating that the $H_{0}$ tension is relieved to $0.47\sigma$, and in this case the model is favored over $\Lambda$CDM ($\Delta{\rm AIC}=-2.07$). In the CBS+$H_{0}$+$S_{8}$ case, we obtain a synthetically best situation, in which $H_{0}=71.9\pm 1.2$ ${\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $S_{8}=0.8074\pm 0.0098$; thus, the $H_{0}$ and $S_{8}$ tensions are relieved to $0.72\sigma$ and $2.11\sigma$, respectively. In this case, the model is most favored by the data ($\Delta{\rm AIC}=-2.29$).
@@
Therefore, we find that such a cosmological model can greatly relieve the $H_{0}$ tension and simultaneously alleviate the $S_{8}$ tension.
@@
Undoubtedly, this model requires further in-depth research in many aspects. Such an interaction between vacuum energy and matter is likely to introduce many additional observational effects. For example, we are not certain whether this model will lead to a significant integrated Sachs-Wolfe effect. Research on this issue requires further related theoretical study and analysis using full CMB angular power spectrum data. Moreover, this additional interaction may also lead to modifications in the middle-scale clustering patterns of large-scale structures, which will potentially affect the clustering strength at various scales. Therefore, such modifications may also affect the mass function and profile of dark matter halos, the statistical properties of galaxy clusters, the alignments of galaxies, the structure of the cosmic web, and so on. The study of these effects is a complicated issue, requires the use of N-body simulations, and also relies on assumptions about the nature of dark matter (such as cold or fuzzy dark matter). All these aspects deserve further in-depth discussion.